Help please!
7 of 10 (>
Answer #1

Given, Pressure P = 9 * 10-13

Radius = 2 * 10-10 m

The mean free path, \lambda = RT / √2 \pid2NAP

R = 8.314, T = 302 K

Na = 6.022 * 1023

The diameter, d = 2 * 2 * 10-10 m

P = 9 * 10-13 * 101325 Pa = 911925 * 10-13

\lambda = 8.314 * 302 / √2 \pi * (2 * 2 * 10-10)2 * 6.022 * 1023 *911925 * 10-13

= 2510.828 / 3.903 * 108 * 10-10

   = 64330.72